Three times a number, minus 6, is equal to two times a number, plus 8. Find the number.
Answer by mananth(16949)

let number be x
3x-6 =2x+8
add +6 to both sides
3x-6+6 = 2x+8+6
3x=2x+14
add -2x
3x-2x= 2x-2x+14
x=14